# Hunt Globally: Deep Research AI Agents for Drug Asset Scouting in Investing, Business Development, and Search & Evaluation

### v1

@@ Abstract
Drug assets scouting is a time-consuming task for many investors and Business Development and Search & Evaluation professionals in the biopharma field. Often, the data is already available on the public web and can be accessed by modern Deep Research AI agents without the need to purchase expensive proprietary software. However, general Deep Research agents still cannot match the quality of human experts in identifying all potential drug assets that meet the search criteria and fit complex queries. In this paper, we propose a robust benchmarking methodology for the drug asset scouting problem and describe our approach of tuning a tree-based self-learning Bioptic Agent to perform complete and non-hallucinated asset scouting. To achieve this, we utilize a multilingual, multi-agent system to construct a challenging completeness benchmark comprising complex user queries and primarily out-of-US-radar regional drugs as the ground truths. To ensure the benchmark reflects real biopharma deal complexity, we collected screening queries from expert investors, BD, and VC professionals and used them as priors to conditionally generate benchmark queries. As graders resolving aliases and conducting research on the extraction of up-to-date attributes for candidate matches, we utilized LLMs as judges, aligned with expert opinions. We compared the performance of our agent against Claude Opus 4.6, OpenAI’s GPT-5.2 Pro, Perplexity’s Deep Research, Gemini 3 Pro + Deep Research, and Exa Websets. Bioptic Agent substantially outperformed existing Deep Research agents, scoring an F1-score of 79.7% versus Claude’s Opus 4.6 (high) (56.2%), Gemini 3 Pro + Deep Research (50.6%), OpenAI’s GPT-5.2 Pro (high) (46.6%), Perplexity’s Deep Research (44.2%), and Exa Websets (26.9%). Our agent also has a steep learning curve, supporting the common belief that spending more compute improves the result.
@@ Introduction
Large pharmaceutical companies rely heavily on external innovation to sustain and expand their pipelines, with the majority of new drugs sourced externally through Business Development and Search & Evaluation (BD and S&E) functions. While pharmaceutical companies have invested significantly in AI to accelerate internal discovery and development, these efforts represent only a small fraction of pipeline development. In contrast, very little progress has been made in applying AI to drug asset scouting, where a substantial share of strategic value lies. In this high-stakes environment, missing a single qualifying program/asset can mean losing a top partnering or acquisition opportunity worth billions of US dollars, making BD and S&E a coverage-critical competition where speed and completeness matter (Schuhmacher et al. 2025; McKinsey & Company 2025).
@@
Crucially, this scouting problem is increasingly global and often disclosed through regional channels, with a substantial amount of innovation coming from outside of the US. For example, WIPO patent applications by origin (origin data) underscore how widely innovation is distributed. U.S.-resident applicants account for ~13.5% of global applications, implying that ~86.5% originate from outside the U.S. (World Intellectual Property Organization 2025). Within this distribution, the largest single origin is China, accounting for ~48.2% of worldwide applications (World Intellectual Property Organization 2025). In biopharma, public leadership commentary has similarly emphasized that major portions of development activity now sit outside traditional U.S./EU hubs. For instance, as reported, Pfizer CEO Albert Bourla cited China as ~30% of global drug development with ~1,200 novel drug candidates (Reuters 2025). Together, these signals motivate a practical shift for scouting where competitive advantage can depend on identifying qualifying programs from heterogeneous, multilingual sources before they are broadly amplified. In such settings, even modest omission rates can be high-cost, which makes completeness and verification central evaluation targets.
@@
As a result, BD and S&E activities are limited to largely manual and time-consuming tasks, relying on human expertise. To maintain a competitive edge and secure top assets, what is needed are deep-research AI agents for global asset scouting: systems that can systematically identify best-in-class assets across languages and sources faster than rivals and optimize for completeness and match or surpass human expert quality.
@@
Existing deep-research tools are typically optimized for fast web fact-finding and citation-backed report synthesis, and common benchmarks reflect this objective. Benchmarks like BrowseComp (Wei et al. 2025) emphasize short browsing tasks with a single verifiable answer, while rubric-based benchmarks such as ResearchRubrics (Sharma et al. 2025) and DRACO (Zhong et al. 2026) evaluate grounding, reasoning, clarity, and citation quality in long-form outputs. Consequently, these evaluation targets favor depth over breadth, where breadth here means completeness-first, open-world set discovery. In BD-style “find-all” searches, this can lead to superficial retrieval and the most novel and valuable programs missed, including those under-the-radar or disclosed primarily in non-English ecosystems. DeepSearchQA moves toward exhaustive answer sets, but it typically evaluates tasks where the correct set is smaller, whereas our setting emphasizes larger-cardinality long-tail enumeration where valid outputs can reach hundreds or thousands of entities (Gupta et al. 2026). Wide-coverage collection benchmarks further show that exhaustive enumeration is a distinct capability that remains difficult even for strong agents in open-world “enumerate-all” settings (Wong et al. 2026), and complementary evaluation work cautions that apparent progress on curated tasks can mask persistent omission and shallow-coverage failure modes in realistic browsing and extraction (Xue et al. 2025). Hence, current Deep Research agents are still less reliable than human experts at consistently identifying all assets that meet complex, multi-constraint criteria.
@@
One of the reasons why this remains unresolved is that much of the required evidence is public but scattered across heterogeneous, fast-moving, and often non-English sources such as company pages, local press, regional trial registries, conference materials, regulatory and legal filings, and corporate PDFs, which forces scouts to traverse many documents to confirm tight criteria while maintaining provenance and completeness under limited time and budget. The same asset frequently appears under multiple aliases through code-name changes, transliterations, and subsidiary disclosures, and early-stage under-disclosure to protect Intellectual Property (IP) is common, so “what exists” and “what can be found reliably” diverge in real workflows. Some notable data vendors such as Clarivate, Global Data have started using LLMs to curate drug assets. However, their databases still lack real-time updates, coverage and cannot handle complex queries. Another point is that the challenge is not only to build an index of all the assets in the world, but also to translate an investor- or diligence-style query into the complete set of matching assets, a step that requires expert interpretation of the technical criteria baked into the query.This intelligence gap is one of the main focuses of our method and evaluation, with precision and coverage as the core metrics.
@@
This paper expands our previous work on benchmarking LLM-based agents for competitive landscape mapping in drug asset due diligence (Vinogradova et al. 2025) by moving from competitor discovery within a single diligence context to open-world, multilingual, “find-all” asset scouting with controlled query intent and difficulty. In a head-to-head evaluation against state-of-the-art commercial deep-research baselines, Bioptic Agent achieves 79.7% F1-score (harmonic mean of precision and recall; higher is better) and substantially outperforms Claude Opus 4.6 high at 56.2%, Gemini 3 Pro with Deep Research at 50.6%, OpenAI GPT-5.2 Pro high at 46.6%, Perplexity Deep Research at 44.2%, and Exa Websets at 26.9%, supporting the conclusion that investor-grade and BD-grade scouting requires completeness-oriented search control, lossless candidate tracking, and expert-aligned validation rather than increased browsing compute or better synthesis alone.
@@
[G1] Find small-molecule antigen presentation modulators in oncology (preclinical through Phase 2). Include only assets with in vivo tumor efficacy and $\leq 3$ more-advanced direct same-target competitors globally. [G2] Find all drug assets targeting LAT1, across any modality, any indication, and any development phase, focusing exclusively on LAT1 (do not include multi-target assets unless LAT1 is clearly a primary target). [G3] Find China-developed/licensed (or have China trial activity) biologic assets (mAbs/bispecifics) in development for autoimmune diseases that modulate the TSHR–IGF-1R axis. [G4] Find all drug assets currently in preclinical or clinical development for treatment of DMD.
@@
To evaluate agent performance on the completeness benchmark, we employ LLM-based graders that assess precision and recall against ground-truth assets, in a similar way we did in (Vinogradova et al. 2025). Both graders use GPT-5.1 with temperature $0.2$ and a fixed seed for reproducibility, and are equipped with web search tools to verify drug attributes and resolve aliases.
@@ Conclusion
In this paper we present (i) a challenging Completeness Benchmark for drug asset scouting and (ii) Bioptic Agent, a tree-based, multilingual deep-research system optimized for complete, non-hallucinated “find-all” scouting. The benchmark is constructed backward from validated program records, mined predominantly from non-U.S., non-English ecosystems, and paired with investor-native, multi-constraint screening queries.
On this benchmark, Bioptic Agent achieves the best overall performance with F1-score of 79.7% and substantially outperforms state-of-the-art commercial deep-research baselines (Table 2), including Claude Opus 4.6 high (0.562), Gemini 3 Pro Deep Research (0.506), o4-mini-deep-research (0.489), GPT-5.2 Pro high (0.466), Perplexity Sonar Deep Research high (0.442), and Exa Websets (0.269). Figure 1 shows a clear quality–time tradeoff: Bioptic Agent improves rapidly early and then approaches a plateau near $\sim 0.80F1$, while sequential “run longer” scaffolds improve more slowly and plateau earlier at lower quality.

### v2

# Hunt Globally: Wide Search AI Agents for Drug Asset Scouting in Investing, Business Development, and Competitive Intelligence

@@ Abstract
Bio-pharmaceutical innovation has shifted: many new drug assets now originate outside the United States and are disclosed primarily via regional, non-English channels. Recent data suggests over 85% of patent filings originate outside the U.S., with China accounting for nearly half of the global total; a growing share of scholarly output is also non-U.S. Industry estimates put China at $\sim 30\%$ of global drug development, spanning 1,200+ novel candidates. In this high-stakes environment, failing to surface "under-the-radar" assets creates multi-billion-dollar risk for investors and business development teams, making asset scouting a coverage-critical competition where speed and completeness drive value. Yet today’s Deep Research AI agents still lag human experts in achieving high-recall discovery across heterogeneous, multilingual sources without hallucinations. We propose a benchmarking methodology for drug asset scouting and a tuned, tree-based self-learning Bioptic Agent aimed at complete, non-hallucinated scouting. We construct a challenging completeness benchmark using a multilingual multi-agent pipeline: complex user queries paired with ground-truth assets that are largely outside U.S.-centric radar. To reflect real deal complexity, we collected screening queries from expert investors, BD, and VC professionals, and used them as priors to conditionally generate benchmark queries. For grading, we use LLM-as-judge evaluation calibrated to expert opinions. On this benchmark, our Bioptic Agent achieves 79.7% F1 score, outperforming Claude Opus 4.6 (56.2%), Gemini 3 Pro + Deep Research (50.6%), OpenAI GPT-5.2 Pro (46.6%), Perplexity Deep Research (44.2%), and Exa Websets (26.9%). Performance improves steeply with additional compute, supporting the view that more compute yields better results.
@@ Introduction
Large pharmaceutical companies rely heavily on external innovation to sustain and expand their pipelines, with the majority of new drugs sourced externally through Business Development and Search & Evaluation (BD and S&E) functions. While pharmaceutical companies have invested significantly in AI to accelerate internal discovery and development, these efforts represent only a small fraction of pipeline development. In contrast, very little progress has been made in applying AI to drug asset scouting, where a substantial share of strategic value lies. In this high-stakes environment, missing a single qualifying program/asset can mean losing a top partnering or acquisition opportunity worth billions of US dollars, making BD and S&E competition where speed and completeness matter [Schuhmacher et al., 2025, McKinsey & Company, 2025].
@@
Crucially, this scouting problem is increasingly global and often disclosed through regional channels, with a substantial amount of innovation coming from outside of the US. For example, WIPO patent applications by origin (origin data) underscore how widely innovation is distributed. Recent data indicates that ~86.5% of global patent applications originate outside the US, with China being the largest country of origin, accounting for ~48.2% of worldwide applications [World Intellectual Property Organization, 2025]. In biopharma, public leadership commentary has similarly emphasized that major portions of development activity now sit outside traditional U.S./EU hubs. For instance, as reported, Pfizer CEO Albert Bourla cited China as ~30% of global drug development with ~1,200 novel drug candidates [Reuters, 2025]. Together, these signals motivate a practical shift for scouting where competitive advantage can depend on identifying qualifying programs from heterogeneous, multilingual sources before they are broadly amplified. In such settings, even modest omission rates can be high-cost, which makes completeness and verification central evaluation targets.
@@
As a result, BD and S&E activities are limited to largely manual and time-consuming tasks, relying on human expertise. To maintain a competitive edge and secure top assets, what is needed are wide-research AI agents for global asset scouting: systems that can systematically identify best-in-class assets across languages and sources faster than rivals and optimize for completeness and match or surpass human expert quality.
@@
Existing deep-research tools are typically optimized for fast web fact-finding and citation-backed report synthesis, and common benchmarks reflect this objective. Benchmarks like BrowseComp [Wei et al., 2025] emphasize short browsing tasks with a single verifiable answer, while rubric-based benchmarks such as ResearchRubrics [Sharma et al., 2025] and DRACO [Zhong et al., 2026] evaluate grounding, reasoning, clarity, and citation quality in long-form outputs. Consequently, these evaluation targets favor depth over breadth, where breadth here means completeness-first, open-world set discovery. In BD-style “find-all” searches, this can lead to superficial retrieval and the most novel and valuable programs being missed, including those under the radar or disclosed primarily in non-English ecosystems. DeepSearchQA moves toward exhaustive answer sets, but it typically evaluates tasks where the correct set is smaller, whereas our setting emphasizes long-tail enumeration with larger cardinality, where valid outputs can reach hundreds or thousands of entities [Gupta et al., 2026]. Wide-coverage collection benchmarks further show that exhaustive enumeration is a distinct capability that remains difficult even for strong agents in open-world “enumerate-all” settings [Wong et al., 2025], and complementary evaluation work cautions that apparent progress on curated tasks can mask persistent omission and shallow-coverage failure modes in realistic browsing and extraction [Xue et al., 2025]. Hence, current Deep Research agents are still less reliable than human experts at consistently identifying all assets that meet complex, multi-constraint criteria.
@@
One of the reasons why this remains unresolved is that much of the required evidence is public but scattered across heterogeneous, fast-moving, and often non-English sources such as company pages, local press, regional trial registries, conference materials, regulatory and legal filings, and corporate PDFs, which forces scouts to traverse many documents to confirm tight criteria while maintaining provenance and completeness under limited time and budget. The same asset frequently appears under multiple aliases through code-name changes, transliterations, and subsidiary disclosures, and early-stage under-disclosure to protect Intellectual Property (IP) is common, so “what exists” and “what can be found reliably” diverge in real workflows. Some notable data vendors, such as Clarivate and GlobalData, have begun using LLMs to curate drug assets. However, their databases still lack real-time updates and coverage, and they cannot handle complex queries. Another point is that the challenge is not only to build an index of all the assets in the world, but also to translate an investor- or diligence-style query into the complete set of matching assets, a step that requires expert interpretation of the technical criteria baked into the query. This intelligence gap is one of the main focuses of our method and evaluation, with precision and coverage as the core metrics.
@@
This paper expands our previous work on benchmarking LLM-based agents for competitive landscape mapping in drug asset due diligence [Vinogradova et al., 2025] by moving from competitor discovery within a single diligence context to open-world, multilingual, “find-all” asset scouting with controlled query intent and difficulty. In a head-to-head evaluation against state-of-the-art commercial deep-research baselines, Bioptic Agent achieves 79.7% F1-score (harmonic mean of precision and recall; higher is better) and substantially outperforms Claude Opus 4.6 high at 56.2%, Gemini 3 Pro with Deep Research at 50.6%, OpenAI GPT-5.2 Pro high at 46.6%, Perplexity Deep Research at 44.2%, and Exa Websets at 26.9%, supporting the conclusion that investor-grade and BD-grade scouting requires completeness-oriented search control, lossless candidate tracking, and expert-aligned validation rather than increased browsing compute or better synthesis alone.
@@
[G1] Find small-molecule antigen presentation modulators in oncology (preclinical through Phase 2). Include only assets with in vivo tumor efficacy and $\leq$ 3 more-advanced direct same-target competitors globally. [G2] Find all drug assets targeting LAT1, across any modality, any indication, and any development phase, focusing exclusively on LAT1 (do not include multi-target assets unless LAT1 is clearly a primary target). [G3] Find China-developed/licensed (or have China trial activity) biologic assets (mAbs/bispecifics) in development for autoimmune diseases that modulate the TSHR–IGF-1R axis. [G4] Find all drug assets currently in preclinical or clinical development for treatment of DMD.
@@
To evaluate agent performance on the completeness benchmark, we employ LLM-based graders that assess precision and recall against ground-truth assets, in a similar way we did in [Vinogradova et al., 2025]. Both graders use GPT-5.1 with temperature $0.2$ and a fixed seed for reproducibility, and are equipped with web search tools to verify drug attributes and resolve aliases.
@@ Conclusion
In this paper we present (i) a challenging Completeness Benchmark for drug asset scouting and (ii) Bioptic Agent, a tree-based, multilingual wide-research system optimized for complete, non-hallucinated “find-all” scouting. The benchmark is constructed backward from validated program records, mined predominantly from non-U.S., non-English ecosystems, and paired with investor-native, multi-constraint screening queries.
On this benchmark, Bioptic Agent achieves the best overall performance with F1-score of 79.7% and substantially outperforms state-of-the-art commercial deep-research baselines (Table 2), including Claude Opus 4.6 high (56.2%), Gemini 3 Pro Deep Research (50.6%), o4-mini-deep-research (48.9%), GPT-5.2 Pro high (46.6%), Perplexity Sonar Deep Research high (44.2%), and Exa Websets (26.9%). Figure 1 shows a clear quality–time tradeoff: Bioptic Agent improves rapidly early and then approaches a plateau near $\sim 0.80$ F1, while sequential “run longer” scaffolds improve more slowly and plateau earlier at lower quality.